# Towards flexible perception with visual memory

### v2

@@ 4 Discussion
Typical neural networks are trained end-to-end: perfect for static worlds, yet cumbersome to update whenever knowledge changes. This is limiting their potential in real-world settings since the world is constantly evolving. Incorporating a visual memory, in contrast, enables a range of flexible capabilities that embrace change: lifelong learning through incorporating novel knowledge, being able to forget, remove and unlearn obsolete knowledge, flexible data selection through memory pruning, and an interpretable decision-making paradigm on which one can intervene to control its behavior. We systematically explored a simple visual memory that decomposes the task of image classification into two primitives, image similarity (from a pre-trained embedding representation) and search (via fast, scalable nearest neighbor search from a vector database). Our results demonstrate that technical improvements like RankVoting improve kNN accuracies for both DinoV2 and CLIP over the widely used SoftmaxVoting method, the previous state-of-the-art, that is sensitive to two hyperparameters (temperature $\tau$ and number of neighbors $k$). Our approach also narrows the accuracy gap between a nearest neighbor memory (best flexibility, perfect unlearning, improved interpretability) and a fixed linear probe (highest accuracy on static image classification). More importantly, we show that visual memory enables flexible perceptual capabilities.
Limitations and future work.
@@
First, we only considered the task of image classification across a broad range of datasets. It will be interesting to extend the approach to other visual tasks, such as object detection, image segmentation, instance recognition and to image generation where a visual memory would also be desirable. Secondly, our approach relies on a fixed, pre-trained embedding model; strong distribution shifts may require updating the embedding. Self-supervised models are a particularly flexible choice, but it is an open question whether one could train smaller models that excel at their task with the help of a larger memory database. Conceptually, if a model needs to save less information in its weights, it might be possible to reduce the computational footprint of such a model. Furthermore, we sometimes observe a trade-off between flexibility and accuracy.
Outlook.
@@
Deep learning is increasingly becoming a victim of its own success: the more widely it is deployed, the stronger its limitations are felt. While the static nature of end-to-end trained networks can easily be forgotten when focusing on fixed, static academic benchmarks, the real world is anything but static. Data is constantly evolving, leading to the dreaded “model drift” where once-optimal models gradually become less effective [69]. Incorporating an explicit visual memory—however it may be instantiated—appears to be a promising way forward, a natural solution for real-world tasks where flexibility is key. While the specific approach we employ here might well be improved through more complex systems, we hope that the flexible capabilities we demonstrated might inspire and contribute to a conversation on how knowledge ought to be represented in deep vision models.
@@ Acknowledgements
We would like to thank Kevin Swersky, David Fleet, Been Kim, Hagen Soltau, Majid Hadian, Kaifeng Chen, Mojtaba Seyedhosseini and Sung Jin Hwang for feedback and discussions, Mike Mozer for helpful comments on our article draft, as well as Rebecca Bland and Phoebe Kirk for legal advice.
[…]
@@ 6. Appendix F
  7. Appendix G
@@
  8. Appendix H
    Algorithm for hierarchical label prediction
  9. Appendix I
@@
  10. Appendix J
    Additional errors on ImageNet-A visualized
  11. Appendix K
@@ Appendix B Aggregation method comparison (iNaturalist)
Figure 8: Aggregating information across retrieved memory samples on iNaturalist. Same as Figure 2 but for iNaturalist instead of ImageNet. (left) Existing aggregation methods (PluralityVoting, DistanceVoting and SoftmaxVoting) are overconfident in distant neighbors, resulting in the paradox of decaying iNaturalist accuracy with more information. (right) This is not the case for our simple RankVoting method which shows strong and stable performance across models and choices of $k$.
@@ Appendix C Hyperparameter sensitivity analysis
Figure 9: Sensitivity to hyperparameters for different aggregation methods. Apart from PluralityVoting, all aggregation methods described in Section 2.2 have a hyperparameter ($\alpha$ for RankVoting, $\tau$ for SoftmaxVoting). For each model and method, we here plot the maximum performance when aggregating using a certain method, sweeping over the number of neighbors from 1 to 100, as a function of the hyperparameter. This analysis is performed to understand how sensitive the respective method is to the choice of the hyperparameter. Note that the x range is different since for instance the temperature parameter in SoftmaxVoting ranges from $[0,1]$ while RankVoting for $\alpha=0$ is undefined (division by zero). We therefore evaluate a broad range for each method and find that all methods have a regime in which they are relatively stable irrespective of the hyperparameter choice. Since DistanceVoting as implemented by Khandelwal et al. [2019] does not have a hyperparameter, we added a temperature-style parameter $\xi$ for the purpose of this comparison by setting $w_{i}=\exp\big{(}-\mathsf{dist}(\tilde{\boldsymbol{z}},\boldsymbol{z}_{[i]})\big{)}^{\xi}$.
Appendix D Hit rate analysis
@@ Appendix E Scaling law
Appendix F NINCO dataset
@@
Figure 11: Distance comparison: the NINCO OOD samples are indeed out-of-distribution for the model. In Section 3.1, we described that we can simply plug new out-of-distribution classes into memory and still perform well on both existing data as well as the new classes. This boxplot confirms that the added samples from the NINCO dataset Bitterwolf et al. [2023] are indeed out-of-distribution for DinoV2 ViT-L14: The mean (left) and median (right) distances from query to the first 100 neighbors are substantially lower for ImageNet validation images than for OOD samples from NINCO.
Appendix G Memory pruning
@@
Appendix H Hierarchical label prediction algorithm for flexibly increasing dataset granularity
Algorithm 1 Hierarchical Label Prediction
1:New example $x$, Hierarchical tree $T$ (with ROOT node)
2:$cur\_node\leftarrow\text{ROOT}$
3:for each level $l$ in $T$ (top to bottom) do
4:     $candidates\leftarrow\text{all\_children\_of}(cur\_node)$
5:     $max\_p\_value\leftarrow-\infty$
6:     $label\_at\_level\leftarrow\text{NULL}$
7:     for each child node $c$ in $candidates$ do
8:         $cross\_dist\leftarrow\text{distance\_distribution}(x,\text{examples}(c))$
9:         $in\_dist\leftarrow\text{distance\_distribution}(\text{examples}(c),\text{examples}(c))$
10:         $p\_value\leftarrow\text{Kolmogorov-Smirnov test}(cross\_dist,in\_dist)$
11:         if $p\_value>max\_p\_value$ then
12:              $max\_p\_value\leftarrow p\_value$
13:              $label\_at\_level\leftarrow c$
14:         end if
15:     end for
16:     $cur\_node\leftarrow label\_at\_level$
17:end for
18:return $label\_at\_level$
Appendix I Calibration analysis
@@
Appendix J Visualizing errors on ImageNet-A
Figure 13: Interpretable decision-making. A retrieval-based visual memory enables a clear visual understanding of why a model makes a certain prediction. Here, we show four randomly selected misclassified query images from ImageNet-A Hendrycks et al. [2021] along with five nearest neighbors from DinoV2 ViT-L14 using the ImageNet-1K training set as visual memory. All labels are from the respective datasets (ImageNet-A for query and ImageNet-train for neighbors). While all neighbors visually look reasonable, not all labels do.
Appendix K Compositionality analysis

### v3

@@ 4 Discussion
Summary. Typical neural networks are trained end-to-end: perfect for static settings, yet cumbersome to update whenever knowledge changes. This is limiting their real-world potential since our world is constantly evolving. Incorporating a visual memory, in contrast, enables a range of flexible capabilities that embrace change: lifelong learning through incorporating novel knowledge, being able to forget, remove and unlearn obsolete knowledge, flexible data selection through memory pruning, and an interpretable decision-making paradigm on which one can intervene to control its behavior. We systematically explored a simple visual memory that decomposes the task of image classification into two primitives, image similarity (from a pre-trained embedding representation) and search (via fast, scalable nearest neighbor search from a vector database). Our results demonstrate that technical improvements like RankVoting improve kNN accuracies for both DinoV2 and CLIP over the widely used SoftmaxVoting method that is sensitive to two hyperparameters (temperature $\tau$ and neighbors $k$). Our approach also narrows the accuracy gap between a nearest neighbor memory (best flexibility, perfect unlearning, improved interpretability) and a fixed linear probe (highest accuracy on static image classification). Importantly, we show that visual memory enables flexible perceptual capabilities.
@@
Limitations and future work. First, we only considered the task of image classification across a broad range of datasets. It will be interesting to extend the approach to other visual tasks, such as object detection, image segmentation, instance recognition and to image generation where a visual memory would be desirable, too (since it is prohibitively expensive to re-train large generative models every time data needs to be removed or added). Secondly, our approach relies on a fixed, pre-trained model; strong distribution shifts may require updating the embedding. Self-supervised models are a particularly flexible choice, but it is an open question whether one could train smaller models that excel at their task with the help of a larger memory database. Conceptually, if a model needs to save less information in its weights, it might be possible to reduce the computational footprint of such a model. Additionally, for scalable vector search, adding/removing samples can require adapting the search index, though the amortized cost is low. Furthermore, we sometimes observe a trade-off between flexibility and accuracy. The use of memory pruning weights as a data selection criterion in the context of dataset pruning (Sorscher et al., 2022) might be an interesting avenue for future work.
@@
Outlook. Deep learning is increasingly becoming a victim of its own success: the more widely it is deployed, the stronger its limitations are felt. While the static nature of end-to-end trained networks can easily be forgotten when focusing on fixed academic benchmarks, the real world is anything but static. Data is constantly evolving, leading to the dreaded “model drift” where once-optimal models gradually become less effective (Bayram et al., 2022). Incorporating an explicit visual memory appears to be a promising way forward for real-world tasks where flexibility is key. While the specific approach we employ here might well be improved through more complex systems, we hope that the flexible capabilities we demonstrated might inspire and contribute to a conversation on how knowledge ought to be represented in vision models.
Impact Statement
The use of a flexible visual memory as proposed in this article has several positive societal implications: (1.) through improved machine unlearning with strong guarantees (by design), it becomes possible for certain classifiers to easily remove data that is no longer considered safe (e.g., complying with user requests or in order to fix data safety issues that are discovered after a model is deployed); (2.) increased transparency and trust through a decision-making mechanism that can be inspected visually and intervened on as opposed to an opaque black-box machine learning system; (3.) potentially increased privacy by storing sensitive data in memory (e.g. on-device) as opposed to training models on the data; (4.) resource efficiency: since fast similarity search can be done on CPUs with a low energy footprint, this suggests possibilities for reducing a system’s carbon footprint (e.g., combining a small model with a larger database).
Given that the system we explore is used for classification, standard classifier-related risks apply, such as for instance the potential for a system to discriminate, amplify biases, or displace human jobs. In addition, the general dual-use problematic of ML models applies: the same system can often be used for beneficial as well as harmful purposes.
We expect that the societal benefits outweigh the harms since the potential harms are shared with all ML classifiers, while the benefits are specific to our approach.
Code availability. Code to replicate experiments from this paper is available at https://github.com/google-deepmind/visual-memory.
@@ Acknowledgements
We would like to thank Kevin Swersky, David Fleet, Been Kim, Hagen Soltau, Majid Hadian, Kaifeng Chen, Mojtaba Seyedhosseini, Ameya Prabhu and Sung Jin Hwang for feedback and discussions, and Mike Mozer for helpful comments on our article draft. We also thank our anonymous reviewers who provided valuable feedback that improved the paper.
@@ 7. Appendix G
  8. Appendix H
    Reliability of retreived neighbors
  9. Appendix I
@@
  10. Appendix J
    Linear probe details
  11. Appendix K
    Latency and storage
  12. Appendix L
    Human experiment to predict model behavior with memory system
  13. Appendix M
@@
  14. Appendix N
    ImageNet-A error analysis
  15. Appendix O
@@
  16. Appendix P
    Connection to bias removal and shortcut learning
@@ Appendix B Aggregation method comparison (iNaturalist)
Figure 8: Aggregating information across retrieved memory samples on iNaturalist. Same as Figure 2 but for iNaturalist instead of ImageNet. (left) Existing aggregation methods (PluralityVoting, DistanceVoting and SoftmaxVoting) are overconfident in distant neighbors, resulting in the paradox of decaying iNaturalist accuracy with more information. (right) This is not the case for RankVoting which shows strong and stable performance across models and choices of $k$.
@@ Appendix C Hyperparameter sensitivity analysis
Figure 9: Sensitivity to hyperparameters for different aggregation methods. Apart from PluralityVoting, all aggregation methods described in Section 2.2 have a hyperparameter ($\alpha$ for RankVoting, $\tau$ for SoftmaxVoting). For each model and method, we here plot the maximum performance when aggregating using a certain method, sweeping over the number of neighbors from 1 to 100, as a function of the hyperparameter. This analysis is performed to understand how sensitive the respective method is to the choice of the hyperparameter. Note that the x range is different since for instance the temperature parameter in SoftmaxVoting ranges from $[0,1]$ while RankVoting for $\alpha=0$ is undefined (division by zero). We therefore evaluate a broad range for each method and find that all methods have a regime in which they are relatively stable irrespective of the hyperparameter choice. Since DistanceVoting as implemented by Khandelwal et al. (2019) does not have a hyperparameter, we added a temperature-style parameter $\xi$ for the purpose of this comparison by setting $w_{i}=\exp\big{(}-\mathsf{dist}(\tilde{\boldsymbol{z}},\boldsymbol{z}_{[i]})\big{)}^{\xi}$.
Appendix D Robustness towards label corruption
Figure 10: Robustness towards label corruption. How robust is a visual memory towards corrupted labels in the memory bank? This plot shows top-1 RankVoting accuracy on the ImageNet validation set as a function of how many labels in the memory (containing ImageNet-1K training set features via DinoV2 ViT-L/14) are corrupted, i.e., assigned to a random class. Intriguingly, performance stays almost unchanged all the way to about 60% (!) corrupted (random) labels in the database.
Appendix E Hit rate analysis
@@ Appendix F Scaling law
Appendix G OOD analysis for NINCO dataset
@@
Figure 12: Distance comparison: the NINCO OOD samples are indeed out-of-distribution for the model. In Section 3.1, we described that we can simply plug new out-of-distribution classes into memory and still perform well on both existing data as well as the new classes. This boxplot confirms that the added samples from the NINCO dataset (Bitterwolf et al., 2023) are indeed out-of-distribution for DinoV2 ViT-L14: The mean (left) and median (right) distances from query to the first 100 neighbors are substantially lower for ImageNet validation images than for OOD samples from NINCO.
Figure 12 confirms that there is a distribution difference between in-distribution data (ImageNet-1K) and OOD data (NINCO). That said, while a distribution shift exists, it is possible that individual NINCO samples were part of the training set for DinoV2. Test-set contamination is generally a concern when working with models trained on large-scale datasets, since test samples may occur as exact, semantic or near-duplicates in large training datasets (e.g. Abbas et al., 2023b). For instance, NINCO contains samples from Food-101 (Bossard et al., 2014) which are also part of LVD-142M dataset used to train DinoV2. That said, the NINCO samples belong to classes which are definitely not part of the ImageNet-train set which serves as a memory bank for our experiments, as ensured by the NINCO dataset collection process (Bitterwolf et al., 2023).
Appendix H Reliability of retrieved memory samples
(a) ImageNet
(b) iNaturalist
Figure 13: Reliability of retrieved memory samples. This plot visualizes the ImageNet (left) and iNaturalist (right) top-1 validation accuracy of a single retrieved neighbor depending on the index of the neighbor (index 0: nearest neighbor). In both datasets and across models, the decrease in accuracy with increasing neighbor index follows smooth trajectories and can be approximated by a two-parameter logarithmic fit (black lines). Figure 13(a) is the same as Figure 1.
Appendix I Memory pruning
@@
Appendix J Linear probe details
For the linear probe results reported in the paper, we directly used the results that were reported in the DinoV2 and CLIP papers. For DinoV2, the authors froze the model backbone and trained the linear layers for 12500 iterations using SGD. Instead of training a single time, they performed a full grid search sweep over three settings (output layers in 1, 4; pooling token concatenation in yes, no, and 13 different learning rates), resulting in 52 linear probes. Then, the authors evaluated the ImageNet validation accuracy for all of those 52 probes and only reported the highest one, as described in Appendix B.3 of the DinoV2 paper. Some may call this test set tuning or double dipping; the DinoV2 paper describes it as “common practice” (Oquab et al., 2023, p. 31). CLIP linear probe results are based on a logistic regression classifier learned using scikit-learn’s L-BFGS implementation, and hyperparameter sweeps are performed on a held-out set not used for evaluation, according to Radford et al. (2021).
Appendix K Latency and storage
Latency.
Nearest neighbor retrieval, fortunately, does not need to reinvent the wheel but can, instead, build on top of highly optimized workloads and libraries such as the ScaNN library (Guo et al., 2020). The ScaNN github README shows a latency comparison; with the requirement of perfect recall a million-size memory can handle roughly 500-600 queries per second. It may be worth mentioning that searching a large database can be done on CPUs and can be heavily parallelized.
Model IN-train features (GB) IN-val features (MB)
DinoV2 ViT-L/14 4.9 197
DinoV2 ViT-B/14 3.7 148
DinoV2 ViT-S/14 1.9 75
CLIP ViT-L/14 3.7 148
CLIP ViT-B/16 2.5 100
Table 10: Storage requirements for ImageNet features. Storing features in a memory database requires only about 1–3% of the space that is needed to store the dataset (154.6 GB for ImageNet-train, 6.0 GB for ImageNet-validation).
Storage.
In addition to latency, storage is another very practical consideration: How much does it take to store features for a large database? To put things into perspective, the ImageNet training dataset requires 154.6 GB of storage, and the ImageNet validation dataset requires 6.0 GB of storage. In comparison, as shown in Table 10, storing DinoV2 or CLIP features for the entire ImageNet training dataset only requires between 1.9 and 4.9 GB of storage space. Thus compared to storing the training dataset, the model features account for only 1–3% of this size. This means that after constructing the memory, one may decide to keep the dataset which adds 1–3% of storage, or one may decide to delete the dataset only keeping the features which saves 97–99% of storage (compared to the dataset storage requirement). The ratio of features requiring 1–3% of the dataset size doesn’t change with dataset scale since it only depends on the embedding model, thus this ratio would hold for very small datasets just as it would for a billion-scale dataset.
Appendix L Human experiment to predict model behavior with memory system
We conducted a small human experiment to quantify how much, if at all, a memory-based system improves interpretability as operationalized by helping humans predict model behavior (as opposed to a standard black-box model). Figure 14 outlines the experimental setup.
Figure 14: Human experiment setup. Conditions A and B were presented on a computer screen in separate trials.
Given 4 label choices (guessing accuracy 25%), human accuracy is 56% in the case of black-box predictions (no neighbor information). With access to four nearest neighbor images from our memory-based system (just the neighbor images but not their labels), human accuracy is at 83%. This represents an absolute improvement of +27% and a relative improvement of +67% in human prediction accuracy, providing strong, falsifiable evidence in favor of the statement that a memory-based model is more interpretable.
Experimental details:
  • The accuracy difference is statistically significant (p ¡ 0.001).
  • Featurizer = DinoV2 ViT-L14 (i.e. the best performing model).
  • Dataset: randomly selected ImageNet-A test images
  • Nearest neighbors for condition B: from ImageNet-train.
  • 4 label choices per trial including ground truth label, model-predicted label (if different), and the remaining 2-3 labels were plausible alternatives based on top CLIP predictions for the test image. Label order randomized.
Appendix M Calibration analysis
@@
Appendix N ImageNet-A error analysis
As shown in Figure 6, many “errors” on ImageNet-A appear to be perfectly reasonable predictions that are caused by dataset label issues as opposed to model mistakes. More randomly selected ImageNet-A samples, along with nearest neighbors, are shown in Figure 16. To quantify the issue, we performed a human experiment on a randomly selected subset of ImageNet-A images (N=100) where the dataset label and the prediction from DinoV2 ViT-L14 with JFT memory disagree. We presented the image alongside the original ImageNet-A label and our model-predicted label to three human observers, asking them to identify which of the labels best describes the image (of course, without telling them which of the labels is the dataset label). The result was that in $39.3\%$ (!) of cases (std: $\pm 1.25\%$), the DinoV2 label was assessed as being better/more suitable than the original dataset label—i.e., roughly 2 out of 5 model “errors” are in fact dataset label errors, quantifying the ImageNet-A label quality issue we alluded to in Figure 6. This percentage can be used to estimate how correcting problematic labels influences performance. Instead of the original model’s $61.1\%$ accuracy on ImageNet-A, due to label errors the ‘corrected’ accuracy is instead $76.4\%$ (a delta of $15.3\%$ in absolute terms or $25.0\%$ in relative terms).
Figure 16: Interpretable decision-making. A retrieval-based visual memory enables a clear visual understanding of why a model makes a certain prediction. Here, we show four randomly selected misclassified query images from ImageNet-A (Hendrycks et al., 2021) along with five nearest neighbors from DinoV2 ViT-L14 using the ImageNet-1K training set as visual memory. All labels are from the respective datasets (ImageNet-A for query and ImageNet-train for neighbors). While all neighbors visually look reasonable, not all labels do.
Appendix O Compositionality analysis
@@
Appendix P Connection to bias removal and shortcut learning
This section contains a brief discussion on a connection between our method and bias removal—since the additional flexibility that our memory approach brings (compared to a standard, inflexible classifier), it is an open question whether this could enable better bias removal.
If the image encoder exploited a shortcut during training, this will influence image similarity and thus nearest neighbor selection. There are cases where bias removal is possible, and cases where it is impossible:
  • Removal impossible: If the encoder is biased towards textures, a test image of “cat shape + elephant texture” (cf. Geirhos et al., 2019, Figure 1) would pull up elephant nearest neighbors, and removing all elephants from memory would come at an unreasonably high cost (not being able to identify elephants anymore). Here, encoder-level debiasing is necessary.
  • Removal possible: If only a part of the memory is biased, memory-level debiasing is feasible. If “fingers” are shortcut predictors for “fish” (due to a dataset bias from proud fishermen holding their catch into the camera, cf. Brendel & Bethge, 2019, Figure 3), then this bias could indeed be rectified by removing the biased “fish+finger” subset from memory. Afterwards, images with “fingers” would no longer lead to “fish” nearest neighbors, demonstrating successful bias removal.